Consider the equation xyz = 4x2 +y2 −z2. Use the Implicit Function Theorem to show that the given equation has a smooth unique local solution of the form z = g(x,y) about the point (2,0,4). Then find the local linearization of g about the point (2,0).


1
Expert's answer
2021-06-09T13:28:44-0400

Assuming "s" is an open subset of "R^{n+k}" and that "F:S\to R^k" is a function of class "C^1" .

Assuming that "(a,b)" is a point in that

"F(a,b)=0" and det "D_yF(a,b)\not =0"

Consider a continuously differentiable function "F(x,y,z)=c" such that

"dF\over dz" "(x,y,z)\not=0" then "F" is "\alpha(x,y,z)" in that "(x,y)" is almost close to "(x,y)" then

"F(x,y,z)=C"

 "xyz=4x^2+y^2-z^2"

"Let F=4x^2+y^2-z^2-xyz"

"\u2207F=[8x\u2212yz,2y\u2212xz,\u22122z\u2212xy]" "\u2207F(2,0,4)=[8\u00d72\u22120\u00d74,2\u00d70\u22122\u00d74,\u22122\u00d74\u22122\u00d70]"

"=[16,\u22128,\u22128]"

Linearization equation

"z=z o\u200b +F x\u200b (x\u2212xy)+F_y\u200b (y\u2212y_1 )"

"z o(x=2,y=0)\space then z=z_o"

"z o\u200b =4\u00d7(2) ^2 +0^2 \u2212z_0^2\n\u200b"

"2_o^2 =16^ 2"

"z_o=4"

"z=4+16(x\u22122)\u22128(y\u22120)"

"=4+16x\u221232\u22128y,"

"16x-8y-z-28=0"